A number is three times the another.If the larger number is subtracted from 60, the result is 5 less than the smaller number subtracted from 55. Find the numbers.
Found 2 solutions by ankor@dixie-net.com, MathTherapy:
Answer by ankor@dixie-net.com(22740)

A number is three times the another.
a = 3b
If the larger number is subtracted from 60, the result is 5 less than the smaller number subtracted from 55.
60 - a = 55 - (b-5)
60 - a = 55 - b + 5
60 - a = 60 - b
This is obviously not going to work, we get
60 - 3b = 60 - b
A bad problem

Answer by MathTherapy(10551)


A number is three times the another.If the larger number is subtracted from 60, the result is 5 less than the smaller number subtracted from 55. Find the numbers.